OpenStudy (caitlinnr14):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (caitlinnr14):

anyone??? do i start with distributive property

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2(x+3) = -2x-6

OpenStudy (anonymous):

that's the left side, the right side is -2x-2-4= -2x-6

OpenStudy (anonymous):

now we have -2x-6=-2x-6 get x's on one side and constants on the other -2x= -2x x=x

OpenStudy (caitlinnr14):

?????

OpenStudy (anonymous):

x=x means "all x" That's the answer.

OpenStudy (caitlinnr14):

ohhhh so all real #numbers

OpenStudy (anonymous):

Right!